# Uncovering Hidden Connections: Iterative Tracking and Reasoning for Video-grounded Dialog

### v1

@@ 3.2.2 Path Tracking
The objective of this phase is to meticulously formulate two discrete entity paths, namely the subject and object paths. These carefully constructed paths encompass entities derived from the dialog history, and these entities bear a substantial resemblance to the corresponding subject and object entities present in the question. This complex process is achieved through two principal sub-processes: intra-utterance interaction and forward tracking, each playing a vital role in the coherence and relevance of the entity paths.
@@
(1) Intra-utterance Interaction. To achieve a robust representation of features, it is imperative to incorporate the co-occurrence information of words, serving as an enabling factor for the subsequent forward tracking step. The refinement of the feature pertaining to the subject in the question is governed by the equation:
@@
where $\mathbf{E}_{tri}[t][0]$, $\mathbf{E}_{tri}[t][1]$, and $\mathbf{E}_{tri}[t][2]$ represent the feature of the subject, relation, and object in the question, respectively. The operation AvgPool($\cdot$) signifies the average pooling mechanism. The term $\overline{\mathbf{E}}_{tri}[t][0]$ captures the enhanced feature of the subject entity embedded within the question. An analogous process yields the enhanced feature for the object entity, epitomized by $\overline{\mathbf{E}}_{tri}[t][2]$. By implementing these computational steps on the triplet details spanning all dialog histories $\mathbf{E}_{tri}\in\mathbb{R}^{t\times 3\times d}$, we derive the refined triplet representation, expressed as $\overline{\mathbf{E}}_{tri}\in\mathbb{R}^{t\times 3\times d}$.
(2) Forward Tracking. In this phase, our focus shifts towards the execution of the forward tracking sub-process, designed to delineate both the subject path and the object path. To offer clarity, we elucidate the mechanism underpinning the establishment of the subject path. For the subject immanent in the question (i.e., the $t$-th utterance), we quantify its similarity with the subject and object entities nestled within the $(t-1)$-th utterance. The decision criterion is encapsulated by the following mathematical formulation:
@@
where $\epsilon_{j}$ represents the similarity score bridging the subject entity in the question and the $j$-th entity in the $(t-1)$-th utterance. The symbol $p$ denotes a predetermined selection threshold, while $S\in\mathbb{R}^{(t-1)\times 2}$ is an adjacency matrix. Note that if both entities in the $(t-1)$-th utterance are larger than $p$, we choose the larger one, as shown in Algorithm 1. The criterion $S[t-1,j/2]=1$ indicates that the $j$-th entity in the $(t-1)$-th utterance resonates with the subject entity in the question. Hence, the aforementioned $j$-th entity is appointed as a pivotal point for the ensuing forward comparison.
Furthermore, it is prudent to evolve from the rudimentary comparison feature, namely the feature of the subject in question $\overline{\mathbf{E}}_{tri}[t][0]$. We propose a more advanced comparison feature, denoted by $\mathbf{h}$, which is structured as:
@@
where $\overline{\mathbf{E}}_{tri}[t][0]$ and $\overline{\mathbf{E}}_{tri}[t-1][j]$ designate the feature of the subject entity in the question and the $j$-th entity (subject or object entity) in the $(t-1)$-th utterance that correlates with the subject entity in the question, respectively. The concatenation operation is symbolized by ||||| |, while $f_{t}$ denotes a fully connected layer, and $\mathbf{h}$ represents the freshly synthesized feature for the ensuing round of comparison.
By perpetuating the forward comparison until reaching the inception of the dialog, a subject path rooted in the subject entity in the question is formulated. A parallel procedure leads to the creation of the corresponding paths, encompassing both the subject and object paths.
Algorithm 1 The path tracking and aggregation mechanism
@@
2:  Initialize adjacency matrix $S$, comparison feature $\mathbf{h}$
@@ 3.2.3 Path Aggregation
In this section, we elaborate on the aggregation mechanism designed to harness the extracted entity paths. The overarching goal is to synthesize a more refined textual representation that encapsulates the crux of the dialog. To achieve this, additional connections are instantiated during the backward traversal of the path, considering the inherent relationships between words. This process has been visually elucidated in Figure 3.
Our strategy leverages an attention mechanism [36], ensuring that during the aggregation process, every node accumulates information from its subsequent nodes, thus refining its representation. For a clearer understanding, let’s consider the case of refining the $i$-th ($i\in\{0,2\}$) entity in the question. Its updated representation is computed as follows:
@@
where $\overline{\mathbf{E}}_{tri}[t][i]$ represents the feature of the $i$-th entity in the $t$-th utterance (the current question), while $\mathcal{N}_{i}$ embodies the set of nodes connected to the $i$-th entity, i.e., all nodes on the same path. Moreover, $\mathbf{E}_{tri}^{txt}[k][j]$ signifies the updated feature with aggregated antecedent information of the $j$-th neighboring entity in the $k$-th utterance. The matrix $\mathbf{W}_{1}$ functions as a parameter matrix used for transformation purposes, and $\mathbf{E}_{tri}^{txt}[t][i]$ connotes the updated feature of the $i$-th entity. An essential aspect of our strategy involves incorporating the distance between utterances into the computation of attention weights. Iteratively applying this mechanism across utterances (from $1$ to $t$) ensures that each entity feature is enhanced. It is worth noting that for nodes with zero in-degree, e.g., entities in the first utterance, we retain their original features.
Following these systematic stages, we derive a richer question representation denoted as $\mathbf{E}_{q}^{txt}\in\mathbb{R}^{n_{q}\times d}$. It is worth noting that $\mathbf{E}_{q}^{txt}$ is constructed by substituting the original word features $\mathbf{E}_{q}$ with the newly refined features from the triplet pairs while leaving word features from other positions intact. Moreover, the word features in the triplet pairs, represented as $\mathbf{E}_{tri}^{txt}[:t-1]$, are regarded as the renewed dialog history representations, symbolized as $\hat{\mathbf{E}}_{h}$. This approach not only curtails the input sequence length but also significantly reduces the perturbations from extraneous words.
@@ 3.3.1 Preprocessing
To achieve a profound extraction of object features from each frame in the video, we resort to the Faster R-CNN model, a renowned convolutional network model optimized for object detection tasks [37, 38, 39]. For our implementation, we draw inspiration from an existing state-of-the-art approach and use a pretrained version of Faster R-CNN [^5] https://github.com/peteanderson80/bottom-up-attention. that was trained on the Visual Genome dataset [14]. Given the richness of videos in terms of the number of frames ($F$) and the plethora of objects ($O$) within these frames, it becomes pertinent to incorporate these features into a singular dimension. Hence, we amalgamate these dimensions, giving rise to the feature representation, $\mathbf{E}^{{}^{\prime}}_{v}\in\mathbb{R}^{{n_{v}}\times{2048}}$. Here, $n_{v}$ is a composite dimension, essentially the product of the frame count and the number of objects, formulated as $n_{v}=F\times O$.
While extraction is a crucial step, the alignment of these extracted features with the textual embeddings (from the question, for instance) is of paramount importance. Features from disparate sources often reside in different dimensional spaces; thus, a mapping mechanism becomes vital to bring congruity between them. For the purpose of achieving this congruence, we integrate a linear layer, further enhanced with the ReLU activation function and layer normalization. This can be mathematically expressed as:
@@
where $f_{v}$ acts as a fully connected layer, functioning as a transformational matrix. The ReLU activation function, symbolized by $\phi$, introduces non-linearity, ensuring the model can learn more complex patterns. Layer normalization, denoted by $\mathrm{Norm}(\cdot)$, scales and shifts the features to ensure they have a mean of zero and a standard deviation of one. The resultant feature set, post this alignment procedure, is represented as $\mathbf{E}_{v}\in\mathbb{R}^{n_{v}\times d}$.
@@ 3.3.2 Iterative Reasoning
For a comprehensive extraction of data related to the question $Q_{t}$ from video $V$, we introduce a novel multimodal iterative reasoning network. This design prioritizes cyclical modality interactions, incrementally refining the data acquired. Both modalities converge into a weight matrix, described as:
@@
where $\mathbf{W}\in\mathbb{R}^{{(n_{q}+n_{v})}\times{(n_{q}+n_{v})}}$ representing the weight matrix and ||||| | indicating concatenation. The matrix is divided into four specific sub-matrices, reflecting different weight components: $\mathbf{W}[:n_{q},:n_{q}]$ for question-question weight, $\mathbf{W}[:n_{q},n_{q}:]$ for question-video weight, $\mathbf{W}[n_{q}:,:n_{q}]$ for video-question weight, and $\mathbf{W}[n_{q}:,n_{q}:]$ for video-video weight.
We utilize four separate attention networks, each capturing distinct semantic information and enhancing the representation across different dimensions. To elucidate, the video-augmented question representations are derived as:
@@
where the weight matrices $\mathbf{W}^{O}$, $\mathbf{W}_{j}^{Q}$, $\mathbf{W}_{j}^{K}$, and $\mathbf{W}_{j}^{V}$ are trainable parameters, and the hyperparameter $d$ scales the dot product within the attention mechanism. The number of attention heads is depicted by $J$, while the video-enhanced question representations are illustrated by $\mathbf{E}_{q}^{v}$. Similarly, the self-enhanced question representations $\mathbf{E}_{q}^{q}$, question-enhanced video representations $\mathbf{E}_{v}^{q}$, and self-enhanced video representations $\mathbf{E}_{v}^{v}$ are obtained through the same procedure.
Subsequently, we amalgamate intra- and inter-modality information to garner more comprehensive representations, as described by,
@@
We recognize that a singular interaction is inadequate to fully elucidate the complex semantic signals within videos. Hence, the described procedure can be cycled, iteratively refining the features. Specifically, $\mathbf{E}_{q}^{vis}$ and $\hat{\mathbf{E}}_{v}$ are considered as new inputs in Eqn. (7) in the next iteration. Multiple iterations appear to foster more efficacious representations for both the question and the video content, thus enhancing the overall system performance. A detailed examination of this iterative process can be found in the Parameter Analysis section.
TABLE I: Performance comparison with the latest methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold and the second in underline.
@@ 3.4 Generator
To deepen our understanding of inquiries, we introduce a neural architecture equipped with gating mechanisms. This architecture adeptly merges relevant information from both dialog history and visual contexts. Specifically, we leverage a dynamic gate regulated by the initial query embeddings $\mathbf{E}_{q}$, text-rich question representations $\mathbf{E}_{q}^{txt}$, and visual-infused question representations $\mathbf{E}_{q}^{vis}$. The formal representation is given as follows:
@@
where $\mathbf{W}_{g}\in\mathbb{R}^{{d}\times{3d}}$ signifies the learnable parameter matrix, the symbol ||||| | represents concatenation, and $\sigma$ is the Sigmoid function. The controlling vector $\mathbf{g}\in\mathbb{R}^{n_{q}\times d}$, comprising elements within the range of 0 and 1, dictates the inclusion of the necessary information, defined as,
@@
where $\odot$ is indicative of the element-wise product operation and $\hat{\mathbf{E}}_{q}$ denotes the ultimate question representation. This gating mechanism flexibly adjusts the influence of each modality during reasoning, ensuring a balanced and holistic query representation.
Building upon the methodology proposed by [15], we utilize the pre-trained GPT-2 model, recognized for its advanced text generation capabilities, as the backbone for our response generator. Specifically, the generator creates responses based on the representations of the video $\hat{\mathbf{E}}_{v}$, the caption $\mathbf{E}_{c}$, the dialog history $\hat{\mathbf{E}}_{h}$, and the posed question $\hat{\mathbf{E}}_{q}$. The combined input for the generator is structured as $\{\hat{\mathbf{E}}_{v},[\mathrm{SEP}],\mathbf{E}_{c},[\mathrm{SEP}],\hat{\mathbf{E}}_{h},[\mathrm{SEP}],\hat{\mathbf{E}}_{q}\}$. Utilizing several transformer layers, the probability of generating a response $A_{t}=\{w_{1}^{a},w_{2}^{a},...,w_{n_{a}}^{a}\}$ is derived as:
$\begin{split}p(A_{t}|V,C,H,Q_{t};\theta)&\approx p(A_{t}|\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta)\\ &=\prod_{i=1}^{n_{a}}p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q}),\end{split}$ (12)
where $n_{a}$ denotes the count of words in the generated response and $\theta$ signifies the parameter of the response generator.
@@ 3.5 Optimization
To refine our optimization strategy, we focus on reducing the negative log-likelihood associated with the actual answer. Employing the Maximum Likelihood Estimation (MLE) criterion facilitates this. The formal expression for the loss is:
$\mathcal{L}_{mle}=-\sum_{i=1}^{n_{a}}\mathrm{log}(p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q})).$ (13)
TABLE II: Details of the benchmark datasets AVSD@DSTC7 and AVSD@DSTC8.
Train Valid Test@DSTC7 Test@DSTC8
Dialogs 7,659 1,787 1,710 1,710
QA pairs 153,180 35,740 13,490 18,810
Words 1,450,754 339,006 110,252 162,226

### v2

@@ 3.2.2 Path Search
The objective of this phase is to construct two discrete entity paths, namely the subject path and object path. These carefully constructed paths encompass entities derived from the dialog history, and these entities bear a substantial resemblance to the corresponding subject and object entities present in the question. This entire process is achieved through two principal sub-processes: intra-utterance interaction and forward search.
@@
(1) Intra-utterance Interaction. To achieve a robust feature representation, it is imperative to incorporate the co-occurrence information of words, serving as a facilitator for the subsequent forward search step. The refinement for the subject entity in the question is governed by the equation:
@@
where $\mathbf{E}_{tri}[t][0]$, $\mathbf{E}_{tri}[t][1]$, and $\mathbf{E}_{tri}[t][2]$ represent the feature of the subject, relation, and object entities in the question, respectively. The AvgPool($\cdot$) denotes the average pooling operation performed on the first axis. The term $\overline{\mathbf{E}}_{tri}[t][0]$ refers to the feature of the subject entity embedded with backbone information in the question. An analogous process yields the enhanced feature for the object entity, denoted as $\overline{\mathbf{E}}_{tri}[t][2]$. By implementing the above steps on the all triplet pairs $\mathbf{E}_{tri}\in\mathbb{R}^{t\times 3\times d}$, we derive the refined triplet representation, expressed as $\overline{\mathbf{E}}_{tri}\in\mathbb{R}^{t\times 3\times d}$.
(2) Forward Search. In this section, our focus shifts towards the execution of the forward search sub-process, which is designed to construct both the subject path and object path. Using the subject path as an example, we illustrate the construction process. For the subject entity in the question (i.e., the $t$-th utterance), we measure its similarity with the subject and object entities within the $(t-1)$-th utterance. The decision criterion is designed by the following mathematical formulation:
@@
where $\epsilon_{j}$ represents the cosine similarity bridging the subject entity in the question and the $j$-th entity in the $(t-1)$-th utterance. The symbol $p$ denotes a predefined selection threshold, while $S\in\mathbb{R}^{(t-1)\times 2}$ is a connection matrix. Note that if both entities in the $(t-1)$-th utterance are larger than $p$, we choose the larger one, as shown in Algorithm 1. The criterion $S[t-1,j/2]=1$ indicates that the $j$-th entity in the $(t-1)$-th utterance resonates with the subject entity in the question. Hence, the aforementioned $j$-th entity is appointed as a pivotal reference point for the ensuing forward comparison.
Furthermore, it is crucial to evolve from the rudimentary comparison feature, namely the feature of the subject in question $\overline{\mathbf{E}}_{tri}[t][0]$. We propose a more advanced comparison feature, denoted by $\mathbf{h}$, which is structured as:
@@
where $\overline{\mathbf{E}}_{tri}[t][0]$ and $\overline{\mathbf{E}}_{tri}[t-1][j]$ are the feature of the subject entity in the question and the $j$-th entity (subject or object entity) in the $(t-1)$-th utterance that correlates with the subject entity in the question, respectively. The concatenation operation is symbolized by ||||| |, while $f_{t}$ denotes a fully connected layer, and $\mathbf{h}$ represents the freshly synthesized feature for the ensuing round of comparison.
By perpetuating the forward comparison until reaching the inception of the dialog, a subject path rooted in the subject entity in the question is formulated. A parallel procedure also leads to the creation of the object path.
Algorithm 1 The path search and aggregation strategy
@@
2:  Initialize connection matrix $S$, comparison feature $\mathbf{h}$
@@ 3.2.3 Path Aggregation
In this section, we elaborate on the aggregation mechanism designed to use the extracted entity paths. The overarching goal is to synthesize a more refined textual representation that integrates the crux of the dialog. To achieve this, additional connections are instantiated during the backward traversal of the path, considering the inherent relationships between words. This process has been shown in Fig. 3.
Our strategy leverages an attention mechanism [39], ensuring that during the aggregation process, every node accumulates information from its antecedent nodes, thus refining its representation. For a clearer understanding, let’s consider the case of refining the $i$-th ($i\in\{0,2\}$) entity in the question. Its updated representation is computed as follows:
@@
where $\overline{\mathbf{E}}_{tri}[t][i]$ represents the feature of the $i$-th entity in the $t$-th utterance (the current question), while $\mathcal{N}_{i}$ embodies the set of nodes connected to the $i$-th entity, i.e., all nodes on the same path. Moreover, $\mathbf{E}_{tri}^{txt}[k][j]$ signifies the updated feature with aggregated antecedent information of the $j$-th neighboring entity in the $k$-th utterance. The matrix $\mathbf{W}_{1}$ functions as a parameter matrix used for transformation and $\mathbf{E}_{tri}^{txt}[t][i]$ connotes the updated feature of the $i$-th entity. An essential aspect of our strategy involves incorporating the distance between utterances into the computation of attention weights. Iteratively applying this mechanism across utterances (from $1$ to $t$) ensures that each entity feature is enhanced. It is worth noting that for nodes with zero in-degree, e.g., entities in the first utterance, we retain their original features.
Following the above process, we derive a richer question representation denoted as $\mathbf{E}_{q}^{txt}\in\mathbb{R}^{n_{q}\times d}$. It is worth noting that $\mathbf{E}_{q}^{txt}$ is constructed by substituting the original word features $\mathbf{E}_{q}$ with the newly refined features from the triplet pairs while leaving word features from other positions intact. Moreover, the word features in the triplet pairs, represented as $\mathbf{E}_{tri}^{txt}[:t-1]$, are regarded as the renewed dialog history representations, symbolized as $\hat{\mathbf{E}}_{h}$. This approach not only shortens the input sequence length but also significantly minimizes disruptions caused by irrelevant words.
@@ 3.3.1 Preprocessing
Following [19, 17], we adopt Faster R-CNN model [^6] https://github.com/peteanderson80/bottom-up-attention. pretrained on the Visual Genome dataset to extract object features of each frame in the video, which is a renowned convolutional network model optimized for object detection task [40, 41, 42]. Given the richness of videos in terms of the number of frames ($F$) and the plethora of objects ($O$) within these frames, we incorporate them into a single dimension, denoted as $\mathbf{E}^{{}^{\prime}}_{v}\in\mathbb{R}^{{n_{v}}\times{2048}}$, where $n_{v}=F\times O$.
Although extraction is a crucial step, aligning these extracted features with the textual embeddings is also paramount. Features from disparate sources often reside in different dimensional spaces; thus, a mapping mechanism becomes vital to bring congruity between them. For this purpose, we integrate a linear layer, further enhanced with the ReLU activation function and layer normalization. This can be mathematically expressed as:
@@
where $f_{v}$ acts as a fully connected layer. The ReLU activation function, symbolized by $\phi$, introduces non-linearity, ensuring the model can learn more complex patterns. Layer normalization, denoted by $\mathrm{Norm}(\cdot)$, scales and shifts the features to ensure they have a mean of zero and a standard deviation of one. Following this alignment procedure, the resulting feature is denoted as $\mathbf{E}_{v}\in\mathbb{R}^{n_{v}\times d}$.
@@ 3.3.2 Iterative Reasoning
For a comprehensive extraction of data related to the question $Q_{t}$ from video $V$, we introduce a novel multimodal iterative reasoning network. This design prioritizes cyclical modality interactions, incrementally refining the information acquired. Both modalities converge into a weight matrix, described as:
@@
where $\mathbf{W}\in\mathbb{R}^{{(n_{q}+n_{v})}\times{(n_{q}+n_{v})}}$ represents the weight matrix, ||||| | indicates concatenation operation, and Softmax($\cdot$) is performed on the last axis. The matrix is divided into four specific sub-matrices, reflecting different weight components: $\mathbf{W}[:n_{q},:n_{q}]$ for question-question weight, $\mathbf{W}[:n_{q},n_{q}:]$ for question-video weight, $\mathbf{W}[n_{q}:,:n_{q}]$ for video-question weight, and $\mathbf{W}[n_{q}:,n_{q}:]$ for video-video weight.
We utilize four separate attention networks, each capturing distinct semantic information and enhancing the representation across different dimensions. For example, the video-augmented question representations are derived as:
@@
where the matrices $\mathbf{W}^{O}$, $\mathbf{W}_{j}^{Q}$, $\mathbf{W}_{j}^{K}$, and $\mathbf{W}_{j}^{V}$ are trainable parameters and the symbol $d$ denotes feature dimension. The number of attention heads is depicted by $J$, while the video-enhanced question representations are illustrated by $\mathbf{E}_{q}^{v}$. Similarly, the self-enhanced question representations $\mathbf{E}_{q}^{q}$, question-enhanced video representations $\mathbf{E}_{v}^{q}$, and self-enhanced video representations $\mathbf{E}_{v}^{v}$ are obtained through the same process.
Subsequently, we integrate intra- and inter-modality information to obtain more comprehensive representations, as described by:
@@
We recognize that a singular interaction is inadequate to fully elucidate the complex semantic signals within videos. Hence, the described procedure can be cycled, $I$ times, iteratively refining the features. Specifically, $\mathbf{E}_{q}^{vis}$ and $\hat{\mathbf{E}}_{v}$ are considered as new inputs in Eqn. (7) in the next iteration. Multiple iterations appear to foster more effective representations for both the question and the video content, thus enhancing the overall system performance. A detailed examination of this iterative process can be found in the Parameter Analysis section.
TABLE I: Performance comparison with the latest methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold and the second is underlined.
@@ 3.4 Generator
To deepen the understanding of inquiries, we introduce a neural architecture equipped with gating mechanisms. This architecture adeptly merges relevant information from both dialog history and visual contexts. Specifically, we leverage a dynamic gate regulated by the initial question embeddings $\mathbf{E}_{q}$, text-rich question representations $\mathbf{E}_{q}^{txt}$, and vision-infused question representations $\mathbf{E}_{q}^{vis}$. The presentation is given as follows:
@@
where $\mathbf{W}_{g}\in\mathbb{R}^{{d}\times{3d}}$ signifies the learnable parameter matrix, the symbol ||||| | represents concatenation, and $\sigma$ is the Sigmoid function. The controlling vector $\mathbf{g}\in\mathbb{R}^{n_{q}\times d}$, comprising elements within the range of 0 and 1, dictates the inclusion of the necessary information, used as follows:
@@
where $\odot$ is indicative of the element-wise product operation and $\hat{\mathbf{E}}_{q}$ denotes the ultimate question representations. This gating mechanism flexibly adjusts the influence of each modality during reasoning, ensuring a balanced and holistic query representation.
Building upon the methodology proposed by [18], we utilize the pre-trained GPT-2 model, recognized for its advanced text generation capabilities, as the backbone for our response generator. Specifically, the generator creates responses based on the representations of the video $\hat{\mathbf{E}}_{v}$, the caption $\mathbf{E}_{c}$, the dialog history $\hat{\mathbf{E}}_{h}$, and the posed question $\hat{\mathbf{E}}_{q}$. The combined input for the generator is structured as $(\hat{\mathbf{E}}_{v},[\mathrm{SEP}],\mathbf{E}_{c},[\mathrm{SEP}],\hat{\mathbf{E}}_{h},[\mathrm{SEP}],\hat{\mathbf{E}}_{q})$. Utilizing several transformer layers, the probability of generating a response $A_{t}=(w_{1}^{a},w_{2}^{a},...,w_{n_{a}}^{a})$ is derived as:
${\begin{split}p(A_{t}|V,C,H,Q_{t};\theta)&\approx p(A_{t}|\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta)\\ &=\prod_{i=1}^{n_{a}}p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta),\end{split}}$ (12)
where $n_{a}$ denotes the word number and $\theta$ signifies the parameter of the response generator.
@@ 3.5 Optimization
To refine our optimization strategy, we focus on reducing the negative log-likelihood associated with the actual answer. Therefore, we employ the Maximum Likelihood Estimation (MLE) [15] loss and adopt the gradient descent method for optimization. The expression for the loss is:
${\mathcal{L}_{mle}=-\sum_{i=1}^{n_{a}}\mathrm{log}\,p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta).}$ (13)
TABLE II: Details of the benchmark datasets AVSD@DSTC7, AVSD@DSTC8, and VSTAR.
Dataset Type Dialogs Turns Words
AVSD@DSTC7 Recorded Video 11K 202K 1.8M
AVSD@DSTC8 Recorded Video 11K 208K 1.9M
VSTAR TV Series 185K 4.6M 30.8M